OpenStudy (anonymous):

Solve: 2x + 36 + 40 = 180 52 128 88 92

OpenStudy (whpalmer4):

Come on, you haven't had any time to try this before asking us.

OpenStudy (whpalmer4):

Subtract 36 and 40 from both sides. What do you get?

OpenStudy (anonymous):

idk ?

OpenStudy (whpalmer4):

What is 180 - 36 -40?

OpenStudy (whpalmer4):

180 -40 ---- ???

OpenStudy (anonymous):

140

OpenStudy (whpalmer4):

140 -36 ---- ???

OpenStudy (anonymous):

104

OpenStudy (whpalmer4):

Right. so the right side of the equation will be 104. Left side:\[2x+36+40-36-40\]What does that equal?

OpenStudy (anonymous):

72?

OpenStudy (whpalmer4):

That's just 2x, right? We had \[2x+36+40 = 180\]We subtracted 36 and 40 from both sides: \[2x+36+40-36-40 = 180-36-40\] \[2x+\cancel{36}+\cancel{40}-\cancel{36}-\cancel{40} = 104\]\[2x=104\]Can you solve that for \(x\)?

OpenStudy (anonymous):

no

OpenStudy (whpalmer4):

No? What number, doubled, gives you 104?

OpenStudy (anonymous):

i dont know

OpenStudy (whpalmer4):

What number is half of 104?

OpenStudy (anonymous):

i dont know omg

OpenStudy (whpalmer4):

Can you divide 104 by 2?

OpenStudy (whpalmer4):

If you had 4 quarters and 2 pennies, you would have 102 cents. How much would half of that be?

OpenStudy (anonymous):

52

OpenStudy (whpalmer4):

Right. \[2x = 104\]Divide both sides by 2 \[\frac{2x}{2} = \frac{104}{2}\]\[\frac{\cancel{2}x}{\cancel{2}} = 52\]\[x=52\]

OpenStudy (anonymous):

so the answee is 52

OpenStudy (whpalmer4):

Yes. Even though I gave you the wrong coin problem :-)

OpenStudy (whpalmer4):

(should have been 4 quarters and 4 pennies = 104, 2 quarters and 2 pennies = 52, but you got the answer we needed)

OpenStudy (texaschic101):

such patience @whpalmer4 ......how do you do it ...lol